OpenStudy (anonymous):

Find the side length of the cube with the volume of 0.343? A. 0.04 cm B. 0.07 cm C. 0.7 cm D. 7 cm

OpenStudy (hoblos):

volume of cube with side s is \[V = s^3\] thus \[s=\sqrt[3]{V}\]

OpenStudy (anonymous):

Okay, thanks but how do i multiply the decimal?

OpenStudy (hoblos):

I don't understand your question clearly

OpenStudy (anonymous):

s = 0.343^(1/3) = 0.7